# Agent: Generating and Evaluating Presentations Beyond Text-to-Slides

### v1

@@ Abstract
Automatically generating presentations from documents is a challenging task that requires balancing content quality, visual design, and structural coherence. Existing methods primarily focus on improving and evaluating the content quality in isolation, often overlooking visual design and structural coherence, which limits their practical applicability. To address these limitations, we propose Agent, which comprehensively improves presentation generation through a two-stage, edit-based approach inspired by human workflows. Agent first analyzes reference presentations to understand their structural patterns and content schemas, then drafts outlines and generates slides through code actions to ensure consistency and alignment. To comprehensively evaluate the quality of generated presentations, we further introduce Eval, an evaluation framework that assesses presentations across three dimensions: Content, Design, and Coherence. Experiments show that Agent significantly outperforms traditional automatic presentation generation methods across all three dimensions. The code and data are available at https://github.com/icip-cas/PPTAgent.
Agent: Generating and Evaluating Presentations Beyond Text-to-Slides
@@ 1 Introduction
Presentations are a widely used medium for information delivery, valued for their visual effectiveness in engaging and communicating with audiences. However, creating high-quality presentations requires a captivating storyline, visually appealing layouts, and rich, impactful content (Fu et al., 2022). Consequently, creating well-rounded presentations requires advanced presentation skills and significant effort. Given the inherent complexity of presentation creation, there is growing interest in automating the presentation generation process (Mondal et al., 2024; Maheshwari et al., 2024) by leveraging the generalization capabilities of large language models (LLM).
@@
Existing approaches often adopt an end-to-end text-generation paradigm, focusing solely on textual content while neglecting layout design and presentation structures, making them impractical for real-world applications. For example, as shown in Figure 1, prior studies (Mondal et al., 2024; Sefid et al., 2021) treat presentation generation as an abstractive summarization task, focus primarily on textual content while overlooking the interactive nature of presentations. This results in simplistic and visually uninspiring outputs that fail to engage audiences.
However, automatically creating visually rich and structurally clear presentations remains challenging due to the complexity of data formats and the lack of effective evaluation frameworks. First, most presentations are saved in PowerPoint’s XML format, which is inherently tedious and redundant (Gryk, 2022). This complex format poses significant challenges for LLMs in interpreting the presentation layout and structure, let alone generating appealing slides in an end-to-end fashion. Second, and more importantly, the absence of comprehensive evaluation frameworks exacerbates this issue. Current metrics like perplexity and ROUGE (Lin, 2004) fail to capture essential aspects of presentation quality such as narrative flow, visual design, and content impact. Moreover, ROUGE-based evaluation tends to reward excessive textual alignment with input documents, undermining the brevity and clarity crucial for effective presentations. These limitations highlight the urgent need for advancements in automated presentation generation, particularly in enhancing visual design and developing comprehensive evaluation frameworks.
Rather than creating complex presentations from scratch in a single pass, presentations are typically created by selecting exemplary slides as references and then summarizing and transferring key content onto them (Duarte, 2010). Inspired by this process, we design Agent to decompose presentation generation into an iterative, edit-based workflow, as illustrated in Figure 2. In the first stage, given a document and a reference presentation, Agent analyzes the reference presentations to extract semantic information, providing the textual description that identifies the purpose and data model of each slide. In the Presentation Generation stage, Agent generates a detailed presentation outline and assigns specific document sections and reference slides to each slide. For instance, the framework selects the opening slide as the reference slide to present meta-information, such as the title and icon. Agent offers a suite of editing action APIs that empower LLMs to dynamically modify the reference slide. By breaking down the process into discrete stages rather than end-to-end generation, this approach ensures consistency, adaptability, and seamless handling of complex formats.
To comprehensively evaluate the quality of generated presentations, we propose Eval, a multidimensional evaluation framework. Inspired by Chen et al. (2024a) and Kwan et al. (2024), Eval leverages the MLLM-as-a-judge paradigm to enable systematic and scalable evaluation. Drawing from Duarte (2010), we categorized presentation quality into three dimensions: Content, Design, and Coherence, providing both quantitative scores and qualitative feedback for each dimension. Our human evaluation studies validated the reliability and effectiveness of Eval.
Results demonstrate that our method effectively generates high-quality presentations, achieving an average score of 3.67 across the three dimensions evaluated by Eval. These results, covering a diverse range of domains, highlight a high success rate of 97.8%, showcasing the versatility and robustness of our approach.
@@
We propose Agent, a novel framework that redefines automatic presentation generation as an edit-based workflow guided by reference presentations.
We introduce Eval, the first comprehensive evaluation framework that assesses presentations across three key dimensions: Content, Design, and Coherence.
We publicly released the Agent and Eval codebase, along with a curated presentation dataset, to facilitate future research in automatic presentation generation.
@@ 2 PPTAgent
In this section, we first establish the formulation of the presentation generation task. Subsequently, we describe the framework of our proposed Agent, which operates in two distinct stages. In stage I, we analyze the reference presentation by clustering similar slides and extracting their content schemas. This process aims to enhance the expressiveness of the reference presentation, thereby facilitating subsequent presentation generation. In stage II, given an input document and the analyzed reference presentation, we aim to select the most suitable slides and generate the target presentation through an interactive editing process based on the selected slides. An overview of our proposed workflow is illustrated in Figure 2.
@@ 2.1 Problem Formulation
Agent is designed to generate an engaging presentation via an edit-based process. We will provide formal definitions for both Agent and the conventional method, illustrating their divergence.
The conventional method for creating each slide can be described in Equation 1, where represents the number of elements on the slide, and denotes the source content composed of sections and figures. Each element on the slide, , is defined by its type, content, and styling attributes, such as .
Compared to the conventional method, Agent adopts an edit-based generation paradigm for creating new slides, addressing challenges in processing spatial relationships and designing styles. This approach generates a sequence of actions to modify existing slides. Within this paradigm, both the input document and the reference presentation serve as inputs. This process can be described as Equation 2, where represents the number of generated actions. Each action represents a line of executable code, and is the reference slide being edited.
@@ 2.2 Stage I: Presentation Analysis
To facilitate presentation generation, we first cluster slides in the reference presentation and extract their content schemas. This structured semantic representation helps LLMs determine which slides to edit and what content to convey in each slide.
@@
Slides can be categorized into two main types based on their functionalities: slides that support the structure of the presentation (e.g., opening slides) and slides that convey specific content (e.g., bullet-point slides). We employ different clustering algorithms to effectively cluster slides in the presentation based on their textual or visual characteristics. For structural slides, we leverage LLMs to infer the functional role of each slide and group them accordingly, as these slides often exhibit distinctive textual features. For the remaining slides, which primarily focus on presenting specific content, we employ a hierarchical clustering approach leveraging image similarity. For each cluster, we infer the layout patterns of each cluster using MLLMs. Further details regarding this method can be found in Appendix C.
@@
After clustering slides to facilitate the selection of slide references, we further analyzed their content schemas to ensure purposeful alignment of the editing. Given the complexity and fragmentation of real-world slides, we utilized the context perception capabilities of LLMs (Chen et al., 2024a) to extract diverse content schemas. Specifically, we defined an extraction framework where each element is represented by its category, modality, and content. Based on this framework, the schema of each slide was extracted through LLMs’ instruction-following and structured output capabilities. Detailed instructions are provided in Appendix E.
@@ 2.3 Stage II: Presentation Generation
In this stage, we begin by generating an outline that specifies the reference slide and relevant content for each slide in the new presentation. For each slide, LLMs iteratively edit the reference slide using interactive executable code actions to complete the generation process.
@@
Following human preferences, we instruct LLMs to create a structured outline composed of multiple entries. Each entry specifies the reference slide, relevant document section indices, as well as the title and description of the new slide. By utilizing the planning and summarizing capabilities of LLMs, we provide both the document and semantic information extracted from the reference presentation to generate a coherent and engaging outline for the new presentation, which subsequently orchestrates the generation process.
@@
Guided by the outline, the slide generation process iteratively edits a reference slide to produce the new slide. To enable precise manipulation of slide elements, we implement five specialized APIs that allow LLMs to edit, remove, and duplicate text elements, as well as edit and remove visual elements. To further enhance the comprehension of slide structure, inspired by Feng et al. (2024) and Tang et al. (2023), we convert slides from their raw XML format into an HTML representation, which is more interpretable for LLMs. For each slide, LLMs receive two types of input: text retrieved from the source document based on section indices, and captions of available images. The new slide content is then generated following the guidance of the content schema.
Subsequently, LLMs leverage the generated content, HTML representation of the reference slide, and API documentation to produce executable editing actions. These actions are executed in a REPL111https://en.wikipedia.org/wiki/Read-eval-print_loop environment, where the system detects errors during execution and provides real-time feedback for self-correction. The self-correction mechanism leverages intermediate results to iteratively refine the editing actions, enhancing the robustness of the generation process.
@@ 3 PPTEval
To address the limitations of existing automated metrics for presentation evaluation, we introduce Eval, a comprehensive framework for assessing presentation quality from multiple perspectives. The framework provides scores on a 1-to-5 scale and offers detailed feedback to guide the improvement of future presentation generation methods. The overall evaluation process is depicted in Figure 3, with the detailed scoring criteria and examples provided in Appendix B.
Drawing from Duarte (2008, 2010), we have identified three key dimensions for evaluating presentation quality:
Content:
Design:
Coherence:
@@ 4.1 Dataset
Data Collection
Existing presentation datasets, such as Mondal et al. (2024); Sefid et al. (2021); Sun et al. (2021); Fu et al. (2022), have two main issues. First, they are mostly stored in PDF or JSON formats, which leads to a loss of semantic information, such as structural relationships and styling attributes of elements. Additionally, these datasets are primarily derived from academic reports, limiting their diversity. To address these limitations, we introduce Zenodo10K, a new dataset sourced from Zenodo (European Organization For Nuclear Research and OpenAIRE, 2013), an open digital repository hosting diverse artifacts from different domains. We have curated 10,448 presentations from this source and made them publicly available to support further research. Following Mondal et al. (2024), we sampled 50 presentations across five domains to serve as reference presentations. Additionally, we collected 50 documents from the same domains to be used as input documents. Details of the sampling criteria are provided in Appendix A.
@@
| Setting | Existing Metrics | PPTEval | ||||||
| Language Model | Vision Model | SR(%) | PPL | FID | Content | Design | Coherence | Avg. |
| Baseline | ||||||||
| GPT-4o | – | – | 110.6 | – | 2.98 | 2.33 | 3.24 | 2.85 |
| Qwen2.5 | – | – | 122.4 | – | 2.96 | 2.37 | 3.28 | 2.87 |
| PPTAgent | ||||||||
| GPT-4o | GPT-4o | 97.8 | 459.7 | 7.48 | 3.25 | 3.24 | 4.39 | 3.62 |
| Qwen2-VL | Qwen2-VL | 43.0 | 322.3 | 7.32 | 3.13 | 3.34 | 4.07 | 3.51 |
| Qwen2.5 | Qwen2-VL | 95.0 | 313.9 | 6.20 | 3.28 | 3.27 | 4.48 | 3.67 |
| Ablation | ||||||||
| PPTAgent | 95.0 | 313.9 | 6.20 | 3.28 | 3.27 | 4.48 | 3.67 | |
| w/o Outline | 91.0 | 2304.3 | 6.94 | 3.24 | 3.30 | 3.36 | 3.30 | |
| w/o Schema | 78.8 | 164.8 | 7.12 | 3.08 | 3.23 | 4.04 | 3.45 | |
| w/o Structure | 92.2 | 189.9 | 7.66 | 3.28 | 3.25 | 3.45 | 3.32 | |
| w/o CodeRender | 74.6 | 231.0 | 7.03 | 3.27 | 3.34 | 4.38 | 3.66 | |
Data Preprocessing
We utilized VikParuchuri (2023) to extract both textual and visual content from the documents. The extracted textual content was then organized into sections using Qwen2.5-72B-Instruct (Yang et al., 2024). For the visual content, captions were generated using Qwen2-VL-72B-Instruct (Wang et al., 2024a). To minimize redundancy, we identified and removed duplicate images if their image embeddings had a cosine similarity score exceeding 0.85. Similarly, slides were excluded if their text embeddings had a cosine similarity score above 0.8 compared to the preceding slide, as suggested by Fu et al. (2022). Detailed statistics of the dataset are presented in Table 1.
4.2 Experimental Settings and Baseline
Models
We evaluate our method using three state-of-the-art models: GPT-4o-2024-08-06 (GPT-4o), Qwen2.5-72B-Instruct (Qwen2.5, Yang et al., 2024), and Qwen2-VL-72B-Instruct (Qwen2-VL, Wang et al., 2024a). These models are categorized according to the specific modalities they handle, whether textual or visual, as indicated by their subscripts. Specifically, we define configurations as combinations of a language model (LM) and a vision model (VM), such as Qwen2.5LM+Qwen2-VLVM.
During experiments, we allow up to two iterations of self-correction per slide generation task, producing presentations per configuration. We use Chen et al. (2024b) and Wu et al. (2020) to compute the text and image embeddings respectively. All open-source LLMs are deployed using the VLLM framework (Kwon et al., 2023) on a cluster of 8 NVIDIA A100 GPUs. The total computational cost for these experiments is approximately 500 GPU hours.
Baseline
We adopt the methodology described in Bandyopadhyay et al. (2024) as our baseline. This approach employs a multi-staged end-to-end model to generate narrative-rich presentations, with an image similarity-based ranking algorithm to add images to the slides. The baseline method is evaluated using either GPT-4o or Qwen2.5, as it does not require the necessary processing of visual information. Each configuration generates presentations, given that it does not require an input presentation. We do not report the success rate and FID of the baseline method for the same reason.

### v3

# PPTAgent: Generating and Evaluating Presentations Beyond Text-to-Slides

@@ Abstract
Automatically generating presentations from documents is a challenging task that requires accommodating content quality, visual appeal, and structural coherence. Existing methods primarily focus on improving and evaluating the content quality in isolation, overlooking visual appeal and structural coherence, which limits their practical applicability. To address these limitations, we propose PPTAgent, which comprehensively improves presentation generation through a two-stage, edit-based approach inspired by human workflows. PPTAgent first analyzes reference presentations to extract slide-level functional types and content schemas, then drafts an outline and iteratively generates editing actions based on selected reference slides to create new slides. To comprehensively evaluate the quality of generated presentations, we further introduce PPTEval, an evaluation framework that assesses presentations across three dimensions: Content, Design, and Coherence. Results demonstrate that PPTAgent significantly outperforms existing automatic presentation generation methods across all three dimensions. The code and data are available at https://github.com/icip-cas/PPTAgent.
PPTAgent: Generating and Evaluating Presentations
Beyond Text-to-Slides
@@ 1 Introduction
Presentations are a widely used medium for information delivery, valued for their visual effectiveness in engaging and communicating with audiences. However, creating high-quality presentations requires a captivating storyline, well-designed layouts, and rich, compelling content (Fu et al., 2022). Consequently, creating well-rounded presentations requires advanced presentation skills and significant effort. Given the inherent complexity of the presentation creation, there is growing interest in automating the presentation generation process (Mondal et al., 2024; Maheshwari et al., 2024; Ge et al., 2025) by leveraging the generalization capabilities of Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs).
@@
Existing approaches typically follow a text-to-slides paradigm, which converts LLM outputs into slides using predefined rules or templates. As shown in Figure 1, prior studies (Mondal et al., 2024; Sefid et al., 2021) tend to treat presentation generation as an abstractive summarization task, focusing primarily on textual content while neglecting the visual-centric nature (Fu et al., 2022) of presentation. This results in text-heavy and monotonous presentations that fail to engage audiences effectively (Barrick et al., 2018).
Rather than creating complex presentations from scratch in a single pass, human workflows typically involve selecting exemplary slides as references and then summarizing and transferring key content onto them (Duarte, 2010). Inspired by this process, we propose PPTAgent, which decomposes slide generation into two phases: selecting the reference slide and editing it step by step. However, achieving such an edit-based approach for presentation generation is challenging. First, due to the layout and modal complexity of presentations, it is difficult for LLMs to directly determine which slides should be referenced. The key challenge lies in enhancing LLMs’ understanding of reference presentations’ structure and content patterns. Second, most presentations are saved in PowerPoint’s XML format, as demonstrated in Figure 11, which is inherently verbose and redundant (Gryk, 2022), making it challenging for LLMs to robustly perform editing operations.
To address these challenges, PPTAgent operates in two stages. Stage I performs a comprehensive analysis of reference presentations to extract functional types and content schemas of slides, facilitating subsequent reference selection and slide generation. Stage II introduces a suite of edit APIs with HTML-rendered representation that simplifies slide modifications through code interaction (Wang et al., 2024b). Furthermore, we implement a self-correction mechanism (Kamoi et al., 2024) that allows LLMs to iteratively refine generated editing actions based on intermediate results and execution feedback, ensuring robust generation. As shown in Figure 2, we first analyze and cluster reference slides into categories (e.g., opening slides, bullet-point slides). For each new slide, PPTAgent selects an appropriate reference slide (e.g., opening slide for the first slide) and generates a series of editing actions (e.g., replace_span) to modify it.
Due to the lack of a comprehensive evaluation framework, we propose PPTEval, which adopts the MLLM-as-a-judge paradigm Chen et al. (2024a) to evaluate presentations across three dimensions: Content, Design, and CoherenceDuarte (2010). Human evaluations validate the reliability and effectiveness of PPTEval. Results demonstrate that PPTAgent generates high-quality presentations, achieving an average score of 3.67 for the three dimensions in PPTEval.
@@
We propose PPTAgent, a framework that redefines automatic presentation generation as an edit-based process guided by reference presentations.
We introduce PPTEval, a comprehensive evaluation framework that assesses presentations across three dimensions: Content, Design, and Coherence.
We release the PPTAgent and PPTEval codebases, along with a new presentation dataset Zenodo10K, to support future research.
@@ 2 PPTAgent
In this section, we formulate the presentation generation task and introduce our proposed PPTAgent framework, which consists of two distinct stages. In stage I, we analyze reference presentations through slide clustering and schema extraction, providing a comprehensive understanding of input presentations that facilitates subsequent reference selection and slide generation. In stage II, we leverage analyzed reference presentations to select reference slides and generate the target presentation for the input document through an iterative editing process. An overview of our workflow is illustrated in Figure 2.
@@ 2.1 Problem Formulation
PPTAgent is designed to generate an engaging presentation through an edit-based process. We provide formal definitions for the conventional method and PPTAgent to highlight their key differences.
The conventional method (Bandyopadhyay et al., 2024; Mondal et al., 2024) for creating each slide is formalized in Equation 1. Given the input content , it generates slide elements, each defined by its type, content, and styling attributes, such as .
While this conventional method is straightforward, it requires manual specification of styling attributes, which is challenging for automated generation (Guo et al., 2023). Instead of creating slides from scratch, PPTAgent generates a sequence of executable actions to edit reference slides, thereby preserving their well-designed layouts and styles. As shown in Equation 2, given the input content and the -th reference slide , which is selected from the reference presentation, PPTAgent generates a sequence of executable actions, where each action corresponds to a line of executable code.
@@ 2.2 Stage I : Presentation Analysis
In this stage, we analyze the reference presentation to guide the reference selection and slide generation. Firstly, we categorize slides based on their structural and layout characteristics through slide clustering. Then, we extract content schemas to identify the content organization of the slide in each cluster, providing a comprehensive description of slide elements.
@@
Slides can be categorized into two main types based on their functionalities: structural slides that support the presentation’s organization (e.g., opening slides) and content slides that convey specific information (e.g., bullet-point slides). To distinguish between these two types, we employ LLMs to segment the presentation accordingly. For structural slides, we leverage LLMs’ long-context capability to analyze all slides in the input presentation, identifying structural slides, labeling their structural roles based on their textual features, and grouping them accordingly. For content slides, we first convert them into images and then apply a hierarchical clustering approach to group similar slide images. Subsequently, we utilize MLLMs to analyze the converted slide images, identifying layout patterns within each cluster. Further details are provided in Appendix D.
@@
After clustering, we further analyzed their content schemas to facilitate the slide generation. Specifically, we define an extraction framework where each element is represented by its category, description, and content. This framework enables a clear and structured representation of each slide. Detailed instructions are provided in Appendix F, with an example of the schema shown below.
Category | Description | Data |
Title | Main title | Sample Library |
Date | Date of the event | 15 February 2018 |
Image | Primary image to illustrate the slide | Picture: Children in a library with … |
@@ 2.3 Stage II : Presentation Generation
PPTAgent first generates an outline specifying reference slides and relevant content for each new slide. Then, it iteratively edits elements from reference slides through edit APIs to create the target presentation.
@@
As shown in Figure 2, we utilize LLM to generate a structured outline consisting of multiple entries. Each entry represents a new slide, containing the reference slide and relevant document content of the new slide. The reference slide is selected based on the slide-level functional description in Stage I, while the relevant document content is identified based on the input document.
@@
Guided by the structured outline, slides are generated iteratively based on the corresponding entries. For each slide, LLMs incorporate textual content and extracted image captions from the input document. The new slide adopts the layout of the reference slide while ensuring consistency in content and structural clarity.
Specifically, to generate a new slide based on the corresponding entry in the outline, we design edit-based APIs to enable LLMs to edit the reference slide. As shown below, these APIs support editing, removing, and duplicating slide elements. Moreover, given the complexity of the XML format in presentations, which is demonstrated in Appendix E, we render the reference slide into an HTML representation (Feng et al., 2024), offering a more precise and intuitive format for easier understanding. This HTML-based format, combined with our edit-based APIs, enables LLMs to perform precise content modifications on reference slides.
Function Name | Description |
del_span | Delete a span. |
del_image | Delete an image element. |
clone_paragraph | Create a duplicate of an existing paragraph. |
replace_span | Replace the content of a span. |
replace_image | Replace the source of image. |
Furthermore, to enhance robustness during the editing process, we implement a self-correction mechanism (Kamoi et al., 2024). Specifically, the generated editing actions are executed within a REPL111https://en.wikipedia.org/wiki/REPL environment. When actions fail to apply to reference slides, the REPL provides execution feedback222https://docs.python.org/3/tutorial/errors.html to assist LLMs in refining their actions. The LLM then analyzes this feedback to adjust its editing actions (Guan et al., 2024; Wang et al., 2024b), enabling iterative refinement until a valid slide is generated or the maximum retry limit is reached.
@@ 3 PPTEval
We introduce PPTEval, a comprehensive framework that evaluates presentation quality from multiple dimensions, addressing the absence of reference-free evaluation for presentations. The framework provides both numeric scores (1-to-5 scale) and detailed rationales to justify each dimension’s assessment.
Grounded in established presentation design principles (Duarte, 2008, 2010), our evaluation framework focuses on three key dimensions, as summarized in Table 1. Specially, given a generated presentation, we assess the content and design at the slide level, while evaluating coherence across the entire presentation.
The complete evaluation process is illustrated in Figure 3, with detailed scoring criteria and representative examples provided in Appendix B.
Dimension | Criteria |
Content | Text should be concise and grammatically sound, supported by relevant images. |
Design | Harmonious colors and proper layout ensure readability, while visual elements like geometric shapes enhance the overall appeal. |
Coherence | Structure develops progressively, incorporating essential background information. |
@@ 4.1 Dataset
Existing presentation datasets, such as Mondal et al. (2024); Sefid et al. (2021); Sun et al. (2021); Fu et al. (2022), have two main issues. First, they are mostly stored in PDF or JSON formats, which leads to a loss of semantic information, such as structural relationships and styling attributes of elements. Additionally, these datasets primarily consist of academic presentations in artificial intelligence, limiting their diversity. To address these limitations, we introduce Zenodo10K, a new dataset sourced from Zenodo (European Organization For Nuclear Research and OpenAIRE, 2013), which hosts diverse artifacts across domains, all under clear licenses. We have curated 10,448 presentations from this source and made them publicly available to support further research.
Following Mondal et al. (2024), we sample 50 presentations in five domains to serve as reference presentations. In addition, we collected 50 documents from the same domains to be used as input documents. The sampling criteria and preprocessing details are provided in Appendix A, while the dataset statistics are summarized in Table 2.
@@
Configuration | Existing Metrics | PPTEval | |||||||
Language Model | Vision Model | SR(%) | PPL | ROUGE-L | FID | Content | Design | Coherence | Avg. |
DocPres (rule-based) | |||||||||
GPT-4o | – | – | 76.42 | 13.28 | – | 2.98 | 2.33 | 3.24 | 2.85 |
Qwen2.5 | – | – | 100.4 | 13.09 | – | 2.96 | 2.37 | 3.28 | 2.87 |
KCTV (template-based) | |||||||||
GPT-4o | – | 80.0 | 68.48 | 10.27 | – | 2.49 | 2.94 | 3.57 | 3.00 |
Qwen2.5 | – | 88.0 | 41.41 | 16.76 | – | 2.55 | 2.95 | 3.36 | 2.95 |
PPTAgent (ours) | |||||||||
GPT-4o | GPT-4o | 97.8 | 721.54 | 10.17 | 7.48 | 3.25 | 3.24 | 4.39 | 3.62 |
Qwen2-VL | Qwen2-VL | 43.0 | 265.08 | 13.03 | 7.32 | 3.13 | 3.34 | 4.07 | 3.51 |
Qwen2.5 | Qwen2-VL | 95.0 | 496.62 | 14.25 | 6.20 | 3.28 | 3.27 | 4.48 | 3.67 |
4.2 Implementation Details
PPTAgent is implemented with three models: GPT-4o-2024-08-06 (GPT-4o), Qwen2.5-72B-Instruct (Qwen2.5, Yang et al., 2024), and Qwen2-VL-72B-Instruct (Qwen2-VL, Wang et al., 2024a). These models are categorized according to the specific modalities they handle, whether textual or visual, as indicated by their subscripts. Specifically, we define configurations as combinations of a language model (LM) and a vision model (VM), such as Qwen2.5+Qwen2-VL.
Experiment data covers 5 domains, each with 10 input documents and 10 reference presentations, totaling 500 presentation generation tasks per configuration (5 domains × 10 input documents × 10 reference presentations). Each slide generation allows a maximum of two self-correction iterations. We use Chen et al. (2024b) and Wu et al. (2020) to compute the text and image embeddings respectively. All open-source LLMs are deployed using the VLLM framework (Kwon et al., 2023) on NVIDIA A100 GPUs. The total computational cost for experiments are approximately 500 GPU hours.
4.3 Baselines
We choose the following baseline methods: DocPres (Bandyopadhyay et al., 2024) propose a rule-based approach that generates narrative-rich slides through multi-stages, and incorporates images through a similarity-based mechanism. KCTV (Cachola et al., 2024) propose a template-based method that creates slides in an intermediate format before converting them into final presentations using predefined templates. The baseline methods operate without vision models since they do not process visual information. Each configuration generates 50 presentations (5 domains × 10 input documents), as they do not require reference presentations. Consequently, the FID metric is excluded from their evaluation.